OpenStudy (anonymous):

At the movie theatre, child admission is $5.10 and adult admission is $9.90 . On Sunday, 141 tickets were sold for a total sales of $1031.10 . How many child tickets were sold that day?

OpenStudy (anonymous):

let's call amount of child admission tickets x and adults - y, then: x+y=141 (total tickets) 5.1*x+9.9y=1,031.10 (total sales) solve this system for x & y

OpenStudy (anonymous):

y=141-x then: 5.1* x +9.9*(141-x)=1031.10 -4.8*x + 1395.9=1031.10 -4.8x=-364.8 x=76 (child admission tickets)

OpenStudy (anonymous):

substitute value of x in original eq & find y: y=141-76 y=65 (adult admission)